OpenStudy (anonymous97):

A ball is thrown from a height of 32 meters with an initial downward velocity of 2/ms . The ball's height h (in meters) after t seconds is given by the following. =h−32−2t5t^2 How long after the ball is thrown does it hit the ground? Round your answer(s) to the nearest hundredth.

OpenStudy (zpupster):

0=−32−2t+5t2 when the ball hits the ground h=0 5t2 -2t - 32 = 0 rearranging solve this with quadratic formula \[-b \pm \frac{ \sqrt { b^2-4ac} }{ 2a }\]

OpenStudy (zpupster):

take the negative answer becasue the balls traveling in a downward direction. Run a check